OpenStudy (anonymous):

Hi, I need help proving the function f:x->x+sin(x) is lipschitz-continuous...

OpenStudy (anonymous):

Definition of a lipschitzian function: \[k \in \mathbb{R}^{+}, \forall (a,b) \in \mathbb{R}^{2}, |f(a)-f(b)|\le k|a-b|\] Now so far I only have that \[|f(a)-f(b)|=|\sin(a)-\sin(b)+a-b|=|2\sin(\frac{a-b}{2})\cos(\frac{a+b}{2})+a-b|\] And I'm not sure about how to use a correct triangular inequality to prove it...

OpenStudy (mr.math):

You can try using the mean value theorem.

OpenStudy (anonymous):

Is it correct to use the Fundamental Theorem of calculus?

OpenStudy (anonymous):

Well, I'm open to any type of solutions, but I can't see how you would use those theorems in this case

OpenStudy (anonymous):

Im still thinking about it, but basically, you want to show that the possible slopes of all the secant lines are bounded.\[|f(a)-f(b)|\leq k|a-b| \iff \left|\frac{f(a)-f(b)}{a-b}\right|\le k\]So if we can show that, we are done.

OpenStudy (anonymous):

You can use the derivative to come to the conclusion that a function is Lipschitz, i dont know if that constitutes as a valid proof though.

OpenStudy (across):

someone1348, doesn't the Lipschitz condition imply that\[|f(a)-f(b)|\leq k|a-b|\implies\left|\frac{f(a)-f(b)}{a-b}\right|\leq k\implies\left|\frac{f(a+h)-f(a)}{h}\right|\leq k\]by putting \(b=a+h\)? This, in turn, would imply\[|f'(x)|\leq k,\]which means all we have do to is determine whether the derivative of \(f\) is bounded or not.

OpenStudy (anonymous):

Yes, that's a geometrical aspect of lipschitz continuity, but I don't know if I can use differentiation, since a and b are not necessarily in a close neighborhood.... And I'd like a general solution that doesn't require that f can be differentiated (my math teacher is a pain, and specifically wanted us to do it without differentiating)

OpenStudy (anonymous):

by using that you only imply that if its lipschitzian then its derivative is bounded, but not the other way around

OpenStudy (zarkon):

you can use the triangle inequality

OpenStudy (anonymous):

That's what I was thinking to use in my first post, but I can't see how to use it right and to set the inequality on the right side :S

OpenStudy (anonymous):

I thought of using, since \[|f(a)-f(b)|=|a-b+2*\sin(\frac{a-b}{2})\cos(\frac{a+b}{2})|\] \[ \leq |a-b|+2|\sin(\frac{a-b}{2})\cos(\frac{a+b}{2})|\iff 2|\sin(\frac{a-b}{2})\cos(\frac{a+b}{2})|\leq (k-1)|a-b|\] But then.... I'm not sure, obviously the left term is bounded, so that means I'm done? Or are there any mistakes?

OpenStudy (anonymous):

hmm, maybe you can prove that f(x) = x and g(x) = sin(x) are individually Lipschitz, then prove that the sum of two Lipschitz functions is also Lipschitz? Let f(x) and g(x) be Lipschitz, and h(x) = f(x)+g(x). Is h(x) also Lipschitz? \[|h(x)-h(y)| = |f(x)+g(x)-f(y)-g(y)|\]\[=|f(x)-f(y)+g(x)-g(y)|\le |f(x)-f(y)|+|g(x)-g(y)|\]because f and g are Lipschitz, we obtain:\[k_1|x-y|+k_2|x-y|=(k_1+k_2)|x-y|\]so we have shown that:\[|h(x)-h(y)|\le K|x-y|\]

OpenStudy (anonymous):

So the sum of two Lipschitz functions is also Lipschitz. Show that f(x) = x and g(x) = sin(x) are Lipschitz.

OpenStudy (jamesj):

This is easy using the triangle inequality: \[ |f(a) - f(b) = | \sin a + a - \sin b - b | \] \[ \leq | \sin a - \sin b| + |a-b| \ \hbox{, by the triangle inequality } \] \[ \leq | a-b| + |a-b| \ \hbox{, by the Mean Value Theorem } \] \[ = 2 |a-b| \]

OpenStudy (anonymous):

....or that lol.

OpenStudy (anonymous):

Thanks! By the way, if \[A=\left\{ k \in \mathbb{R}^{+}, f k-lipschitzian \right\} \] Then \[\inf(A)=1\] Because, by calling B and C the sets of lipschitz constants for sin and x->x, inf(A)=inf(B)+inf(C)=1, no?

OpenStudy (jamesj):

If you're asking whether the inf of the set of Lipschitz constants for f(x) = sin x + x is 1, I'd say no; it's more likely to be 2. Just look at \[ | f(δ)−f(−δ) | \] for arbitrarily small δ

OpenStudy (anonymous):

true... I think, in fact, sin is 1-lipschitzian? Let's see... \[|\sin(a)-\sin(b)|=2|\sin(\frac{a-b}{2})\cos(\frac{a+b}{2})|\] and then I'm stuck =/

OpenStudy (jamesj):

For sure the best bound k such that \[ | \sin x - \sin y | \leq k |x-y| \] is k = 1; just look at it around the origin. To formally prove it: suppose there exists a Lipschitz constant k where k < 1. Then for all x,y \[ \left| \frac{ \sin x - \sin y }{x -y} \right| \leq k < 1 \] in which case the maximum value of the derivative of sin would be k < 1. But this is false. Hence k must be at least 1. Now we can prove with the MVT that it can be 1. Hence the inf(Lip constants of f) = min(Lip constants of f) = 1.

OpenStudy (anonymous):

But to prove it's k=1, you have to first admit that sin is lipschitzian?

OpenStudy (jamesj):

No. use the MVT. This is a very useful result (the result of applying the MVT to the function sin) to figure out and learn if you don't know it.

OpenStudy (jamesj):

If you mean we need to show first sin is Lipschitz, yes.

OpenStudy (jamesj):

...and the way to do it is with the MVT.

OpenStudy (anonymous):

Is there any way to do it without it?

OpenStudy (jamesj):

Why? There's no reason not to use a result from first year calculus in a course in analysis, which is where Lipschitz continuity belongs.

OpenStudy (anonymous):

Forgot to state my problem asks first to prove f is lipschitzian, then to calculate inf(a) without using the mean value theorem... I forgot to say it because the problem is in french, and I didn't know that the theorem I knew was MVT in english... :S And yeah... but she specifically prohibited us from using it in this problem, which is why I think more and more that french math is twisted...

OpenStudy (zarkon):

you can use the unit circle and a simple geometric argument to show that \[\sin(\theta)\le \theta\]

OpenStudy (zarkon):

or better \[|\sin(\theta)|\le |\theta|\]

OpenStudy (anonymous):

Hmm true, especially since before they had me prove that \[|\sin(x)|\leq x\]

OpenStudy (anonymous):

Yep, I have that, it was a particular pain to write the proof formally, but then afterwards what do you suggest I do with it?

OpenStudy (zarkon):

use it on \[|\sin(a)-\sin(b)|=2|\sin(\frac{a-b}{2})\cos(\frac{a+b}{2})|\]

OpenStudy (jamesj):

Then use the identities you've been playing with above: \[ | \sin x - \sin y | = 2 | \sin((x-y)/2) . \cos((x+y)/2) | \] \[ \leq ... what? \]

OpenStudy (jamesj):

You know you want |x-y| on the right-hand side. See how to get there.

OpenStudy (anonymous):

\[\leq|x-y|\cos(\frac{x+y}{2})\] ?

OpenStudy (zarkon):

now bound the cosine function

OpenStudy (jamesj):

yes and someone doing analysis should know how to bound cos

OpenStudy (anonymous):

...yep, duh. crap. I feel extremely bad now hahaha

OpenStudy (anonymous):

Thanks to you both.